Use the information in the table below to answer questions 1 and 2:

GDP at market prices R378bn

Net primary income payments to the rest of the world R37 bn

Indirect taxes R51bn

Subsidies R12 bn

Consumption of fixed capital R43 bn



GDP at factor cost is R339 bn. TRUE OR FALSE?


2) The value of Gross National Income (GNI) at market prices is:R341 bn. TRUE OR FALSE?
1
Expert's answer
2021-09-03T13:50:25-0400
"solution"

"GDP_{fc}=GDP_{mp }\ -Indirect\ taxes+subsidies\\
339=378-51+12\\
339=390-51\\
339=339.\\"


GDP at factor cost is R339bn therefore TRUE.


"GNI_{mp}=GDP_{mp}+NFIA(EX_{fs}-IM_{fs})\\
341=378+(0-37)\\
341=378-37\\
341=341"


Gross National Income (GNI) at market prices is:R341 bn. TRUE